OpenStudy (anonymous):

H-E-L-P Determine the extrema of below on the given interval f(x)=5x^3-61x^2+16x+3 (a) on [0,4] The minimum is ?? and the maximum is ?? (b) on [-9,9] The minimum is ?? and the maximum is ?? please solve so i can know which answer i got wrong.

OpenStudy (anonymous):

Do you know the derivative?

OpenStudy (anonymous):

yes i did everything. x= 2/15 and x= 8

OpenStudy (anonymous):

those are the critical points.

OpenStudy (anonymous):

left end point: f(0)= 3 critical points: x= 2/15 f(2/15)= 4.06074 x=8 f(8)= -1213 Right Point: f(4)= -589

OpenStudy (anonymous):

Okay you don't consider \(f(8)\) for the \([0,4]\) interval..

OpenStudy (anonymous):

oh......

OpenStudy (anonymous):

It's apparent that \(f(2/15)\) is the maximum.

OpenStudy (anonymous):

And the minimum is going to be \(f(4)\)

OpenStudy (anonymous):

just 2/15? or do we write that the max as: 4.06074

OpenStudy (anonymous):

so the min has to be: -589

OpenStudy (anonymous):

I said \(f(2/15)\) which is equal to \(4.06074\)

OpenStudy (anonymous):

okay. then what's the minimum

OpenStudy (anonymous):

Now you just need to check 9 and -9 for the other interval.

OpenStudy (anonymous):

since i can't use f(8)

OpenStudy (anonymous):

For the \([0,4]\) interval, the minimum is -589 right?

OpenStudy (anonymous):

i think so...

OpenStudy (anonymous):

For the \([-9,9]\) interval... you can consider \(f(8)\), but you also need to consider \(f(-9)\) and \(f(9)\)

OpenStudy (anonymous):

after i do all the work, the problem is i don't know how to pick the min and max.

OpenStudy (anonymous):

for f(-9)= -8727 and for f(9) = -1149.

OpenStudy (anonymous):

Is this correct ?? (a) on [0,4] The minimum is -589 and the maximum is 4.06074 (b) on [-9,9] The minimum is -8727 and the maximum is -1149

OpenStudy (anonymous):

Yeah... the minimum is the LOWEST value the maximum is the HIGHEST value.

OpenStudy (anonymous):

You only have to consider the values at critical points and endpoints.

OpenStudy (anonymous):

i submitted this as my answer but it says its wrong. :(

OpenStudy (anonymous):

You must be doing it wrong then... somehow.

OpenStudy (anonymous):

i know that's why im asking you if the answer i had up there are correct.

OpenStudy (anonymous):

What's your derivative?

OpenStudy (anonymous):

15x^2-122x+16

OpenStudy (anonymous):

can you get back to me with answer so i can see if its correct. i dont want to go step by step if its wrong in the end. i tried this problem several times...